Income is $99,000, paid $18,900 in taxes, what fraction shows how much spent on taxes?

Answer by stanbon(75887)

Income is $99,000, paid $18,900 in taxes, what fraction shows how much spent on taxes?
------
Ans:: 18900/99000 = 189/990 = 63/330 = 21/110
------------
Cheers,
Stan H.
-------------